A 0.14-M solution of a weak acid is 3.8% dissociated. Calculate K. K =
Answer #1

% dissociation = (x*100)/c

3.8= x*100/0.14

x = 5.32*10^-3

HA dissociates as:

HA -----> H+ + A-

0.14 0 0

0.14-x x x

Ka = [H+][A-]/[HA]

Ka = x*x/(c-x)

Ka = 5.32*10^-3*5.32*10^-3/(0.14-5.32*10^-3)

Ka = 2.101*10^-4

Answer: 2.10*10^-4